OpenStudy (anonymous):

Two numbers which sum = 28 and product= -60 will medal!

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@AdamK

OpenStudy (adamk):

You know it's going to be a positive number and an negative number since when they are multiplied you get a negative number.

OpenStudy (adamk):

And you know the absolute value of the positive number will be greater than the absolute value of the negative number since it's a positive number when added.

OpenStudy (anonymous):

yes I know, ive been working on these two questions and Im stuck

OpenStudy (adamk):

Ok, there actually aren't two numbers that add to get 28 and multiply to get -60.

OpenStudy (anonymous):

That's what my teacher is askng

OpenStudy (anonymous):

the other one is sum of 56 and product of 300

OpenStudy (adamk):

The first one is 30 and -2

jimthompson5910 (jim_thompson5910):

Here are all of the possible ways to multiply 2 whole numbers to get -60 1*(-60) = -60 2*(-30) = -60 3*(-20) = -60 4*(-15) = -60 5*(-12) = -60 6*(-10) = -60 (-1)*(60) = -60 (-2)*(30) = -60 (-3)*(20) = -60 (-4)*(15) = -60 (-5)*(12) = -60 (-6)*(10) = -60 You then make a table of all the sums of the pairs (see attached)

OpenStudy (anonymous):

thanks @AdamK an I have help on the second one

OpenStudy (anonymous):

please :(